# IAFormer: Interaction-Aware Transformer network for collider data analysis

### v1

@@ Abstract
In this paper, we introduce IAFormer, a novel Transformer-based architecture that efficiently integrates pairwise particle interactions through a dynamic sparse attention mechanism. The IAformer has two new mechanisms within the model. First, the attention matrix depends on predefined boost invariant pairwise quantities, reducing the network parameter significantly from the original particle transformer models. Second, IAformer incorporate the sparse attention mechanism by utilizing the “differential attention”, so that it can dynamically prioritizes relevant particle tokens while reducing computational overhead associated with less informative ones. This approach significantly lowers the model complexity without compromising performance. Despite being computationally efficient by more than an order of magnitude than the Particle Transformer network, IAFormer achieves state-of-the-art performance in classification tasks on the Top and quark-gluon datasets. Furthermore, we employ AI interpretability techniques, verifying that the model effectively captures physically meaningful information layer by layer through its sparse attention mechanism, building an efficient network output that is resistant to statistical fluctuations. IAformer highlights the need to sparse attention in any Transformer analysis to reduce the network size while improving its performance.
@@ 1 Introduction
Analyzing LHC collision data to uncover intriguing physics phenomena, selecting an appropriate jet reconstruction approach and tagging algorithm is essential, as they must align with the expected characteristics of the final state. For scenarios involving the production of a high momentum, hadronically decaying heavy particle such as a W, Z, or a top quark, a large radius jet is often used to capture entire particles arising from the heavy particle decay.
The recognition of fat jets with the multi-prong structure was first noted in Ref. [1], where jet substructure techniques emerged as a powerful method for suppressing QCD backgrounds, thereby advancing Beyond Standard Model (BSM) searches at the LHC. The pronged structure associated with different heavy resonant particles is dictated by the final state products of their decays. Over time, jet substructure methodologies have evolved, leading to Deep Learning (DL) based classification techniques for tagging heavy particles [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. These approaches are now widely applied in LHC analyses conducted by ATLAS and CMS [22, 23, 24, 25, 26, 27]. They are also employed in scenarios with multiple heavy particle final states [28, 29, 30] and in extracting fundamental parameters of the theory [31, 32, 33, 34, 35]. In contrast, processes that generate spatially isolated quarks or gluons are typically reconstructed using one or more small radius jets. This is followed by flavor tagging, a technique that identifies the type of initiating parton, whether an up, down, strange, charm, or bottom quark, or a gluon. Bottom and charm hadrons (primarily B and D hadrons) have relatively long lifetimes, allowing them to travel a short distance within the detector before decaying. Their decays produce displaced tracks originating from a secondary vertex, distinct from the primary collision point. By leveraging machine-learning techniques, this vertex displacement information, combined with subsequent decay properties, significantly improves tagging efficiency [36, 37, 38, 39]. Additionally, studies have explored the potential for strange-quark identification using ML techniques [40, 41, 42]. Since gluons belong to the octet representation of QCD SU(3) symmetry, their jet characteristics differ from those of quarks, which are in the triplet representation. Consequently, ML-based methods have also been developed to distinguish gluon-initiated jets from quark-initiated ones [43, 44, 45, 46, 47, 48].
@@
Recently, Transformer networks have demonstrated significant improvements in jet tagging tasks [49, 50, 21]. In their standard architecture, Transformers process jet constituents along with their corresponding features, which include kinematic information for each individual hadron inside the jet. In this setup, the network assigns weights to each particle based on its feature importance in the overall classification decision. However, incorporating features derived from particle-pair interactions can further enhance network performance, as these features play a crucial role in capturing the radiation pattern inside the jet [10]. Efforts to integrate pairwise interactions into the attention mechanism have involved either using an interaction matrix as a bias for the computed attention scores [49] or replacing the Query and Key matrices with the interaction matrix itself [50]. In both approaches, the number of attention heads is constrained by the number of features in the interaction matrix. Additionally, the interaction matrix remains fixed across attention heads and is injected into each attention layer without updates. This rigid structure can disrupt the learning of attention patterns across different layers, potentially limiting the network overall effectiveness. IAFormer addresses these challenges by allowing the interaction matrix to be updated across different attention heads and Transformer layers. Furthermore, the interaction matrix is refined through skip connections and propagated to later attention layers.
@@
The IAFormer architecture demonstrates superior performance compared to previously introduced Transformer-based models when evaluated on top tagging and quark-gluon tagging tasks, achieving this with a significantly reduced network size. Although the network is validated against state-of-the-art jet tagging models, IAFormer is designed as a general framework that can be adapted for various classification tasks. The implementation is publicly available at IAFormer GitHub, and detailed usage instructions are provided in the appendix.
This paper is organized as follows; Section 2 discusses various Transformer network architectures, including the structure of IAFormer. Section 3 presents validation results, comparing IAFormer with existing networks for top tagging and quark-gluon tagging tasks. In Section 4, we employ different interpretability methods to analyze the learned attention patterns in IAFormer during the top tagging task. Section 5 provides a discussion and conclusion.
@@ 2 Transformers for jet tagging
Transformers were originally introduced for natural language processing, particularly for sequence-to-sequence tasks like machine translation [52], which were later modified for collider physics. At the core of the Transformer is the self-attention mechanism, which dynamically assigns importance to different input components. One of the most significant advantages of self-attention is its ability to capture long-range dependencies, as it considers all elements of the input simultaneously rather than relying on sequential processing. Moreover, modifications to the self-attention formula by incorporating the particle pair interactions enhance the network’s performance.
In jet tagging tasks, the outcome of an event should remain unchanged regardless of the order in which particles are processed. Therefore, the designed DL networks should treat particles as an unordered set, commonly referred to as a “particle cloud”. By representing hardons inside the jet in a permutation-invariant manner, particle cloud models overcome the challenges posed by combinatorial ambiguities. The particle-based methods also provide the flexibility to incorporate additional particle attributes, such as particle identification and vertex information.
Several machine learning architectures have been developed for jet tagging under the permutation invariant framework, including Deep Sets [53], Edge Convolution Networks [54], and Transformers [49, 29, 55, 56, 16, 20]. Deep Sets can achieve state-of-the-art performance but require a large latent space, making them computationally demanding. Edge Convolution Neural Networks (EDGCNN) address this challenge by leveraging local neighborhood information within the particle cloud to enhance learning. For the case of particle transformers, attention weights are computed for every particle interaction within the dataset and outputs are aggregated to keep the permutation information, so that the network output remains unaffected by the order of the input. However, self-attention comes with computational challenges, particularly its quadratic scaling with input size. Various improvements, such as sparse attention [57], have been proposed to address this issue.
In the following, we first introduce the Plain Transformer and its modified version including the interaction matrix, then present the new network, IAFormer. IAFormer significantly reduces the computational burden of attention matrix calculation by incorporating prefixed pairwise information with trainable weights. Moreover, it reduces statistical fluctuations by introducing a dynamic spatial attention mechanism, referred to as differential attention.
@@ 2.1 Plain Transformer
The core component of a transformer model is the attention matrix, which is a weighted sum of the input elements. The weights are determined based on the relevance of each element to the others. By ensuring that the attention weights are computed in a permutation-invariant manner, the resulting model is able to effectively process sets.
Mathematically, the attention mechanism operates as follows: The input to the multi-head attention $X$ has an unordered grid-like structure with dimensions, $I\times J$, with $I$ indicating the number of particles and $J$ indicating the relevant features for each particle in the dataset.
For given $X_{ij}$, the feature vector $X_{*j}$ is embedded to a higher dimension via an MLP. The embedding to a higher dimension helps the network to capture meaningful complex patterns and nonlinear relationships that may not be apparent in the original feature space; therefore, it enhances the model’s ability to distinguish different classes for deep learning tasks like classification and anomaly detection. High-dimensional embeddings also provide richer representations that facilitate downstream tasks such as contrastive learning and clustering. The size of the embedded features is determined by the number of neurons in the final MLP layer as
$X^{\rm embed}_{i,j}=\sigma\left(X^{(l-1)}_{i,j^{\prime}}W^{l}_{j^{\prime},j}+b^{l}_{i}\right)\,,$ (1)
@@
Once the input features are mapped to a higher dimension, the inputs are passed to three separate linear layers forming the Query, Key, and Value matrices as
$Q^{I\times I}_{i,j}\equiv X^{\rm embed}_{i,j^{\prime}}\cdot W^{Q}_{j^{\prime},j}\,,\hskip 11.38109ptK^{I\times J}_{i,j}\equiv X^{\rm embed}_{i,j^{\prime}}\cdot W^{K}_{j^{\prime},j}\,,\hskip 11.38109ptV^{I\times J}_{i,j}\equiv X^{\rm embed}_{i,j^{\prime}}\cdot W^{V}_{j^{\prime},j}\,,$ (2)
where $Q$, $K$ and $V$ are the query, key, and value matrices, respectively, used to compute the attention for the dataset. The suffix $I\times J$ indicates the dimension of the matrices, where $I$ is the number of particle tokens and $J$ is the number of features. The weight matrices have the dimensions of $I\times I$ to preserve the dimensions of the embedded input dataset. The attention score $\alpha$ is defined as:
@@
where in the last line sum runs over all components $I\times I$, suffix $I\times I$ of $\alpha$ indicates the dimension of the attention weight matrix $\alpha$, and $d$ is the length of the particle tokens.
@@ 2.2 Transformer with interaction matrix
[…]
@@
with ${\mathcal{I}}_{i,i}$ is the pairwise interaction matrix for a single head after the embedding. In this structure, the interaction matrix acts as a bias in the attention score. While this approach effectively integrates the interaction matrix into the attention score, it has certain limitations. First, the number of attention heads has to be increased with the number of features of the interaction matrix. If the interaction matrix has a large number of features, the number of attention heads increases accordingly, leading to a significant rise in network parameters. This occurs because $Q$ and $K$ must accommodate the same number of attention heads in this setup. Second, all Transformer layers use the same fixed interaction matrix as a bias, preventing the model from learning updated representations in deeper layers, leading to lower performance.
@@
While this approach reduces network complexity and improves performance, it has a key limitation. Since the interaction matrix is independently added to each Transformer layer. This leads to the mismatch in feature representations between the pairwise interaction matrix and the input $X_{i,j}$. As the attention mechanism assigns weights to all particle tokens in the dataset, this misalignment can lead to assigning attention to irrelevant particles, such as soft hadrons.
@@ 2.3 IAFormer
IAFormer is a transformer-based architecture, but significant modifications have been made compared to the ParT setup, as illustrated in Figure 2.3.2. Instead of computing the attention score through the multiplication of the Query and Key matrices, IAFormer applies the softmax function directly to a trainable interaction matrix of the form $W{\mathcal{I}}$. This interaction matrix is independently optimized for each attention head within the individual attention layers, enabling the network to dynamically learn discriminative patterns among different jets with effectively reduced size.
@@
Additionally, a dynamic sparse attention pattern is included following the idea of “differential attention” [51]. The differential attention utilises the attention scores as the difference between two separate softmax attention maps. The subtraction cancels noise, promoting the sparse attention patterns dynamically. The model has a trainable parameter $\beta$ to control the level of the subtraction. The parameter is initialized randomly and shared across all attention heads within each attention layer. This mechanism suppresses attention scores for less relevant particle tokens, enhancing the model’s focus on important interactions.
@@ 2.3.1 Dynamic sparse attention via differential attention
Sparse attention is a technique used in Transformer models to reduce the computational complexity of the self-attention mechanism by limiting the number of key-query pairs attend to the classification. Instead of computing attention over all tokens in a sequence, sparse attention selectively attends to a subset of tokens based on predefined patterns, learned structures, or efficient approximations. This significantly lowers the memory and computational cost from quadratic to linear or sub-quadratic complexity, making it possible to process longer sequences efficiently.
Different types of sparse attention have been introduced such as fixed pattern and learnable sparse attention. Fixed pattern sparse attention is a type of sparse attention mechanism where each token attends only to a predefined subset of tokens based on a fixed rule.[58, 59, 60]. On the other hand, Learnable sparse attention dynamically determines which tokens to attend to based on learned patterns. Learnable sparse attention allows the model to adaptively focus on the most relevant tokens in different contexts, leading to improved efficiency. In this paper we utilize a dynamic sparse attention mechanism incorporated in transformer called “differential attention” which was first introduced in [51].
@@
with $\beta$ is a learnable vector and $W_{1},W_{2}$ are two learnable matrices. During the training process, the value of $\beta$ vector is optimized to demote the attention score assigned to less relevant tokens, resulting in the attention score matrix to be sparse. This results in an implicit sparsity, where tokens gradually focus only on meaningful interactions while ignoring less relevant ones. The self-attention mechanism is guided by this dynamic sparsity, reducing computational overhead while preserving essential long-range dependencies. The IAFormer shows better performance compared with ParT with notably smaller number of parameters.

### v2

@@ Abstract
In this paper, we introduce IAFormer, a novel Transformer-based architecture that efficiently integrates pairwise particle interactions through a dynamic sparse attention mechanism. IAFormer has two new mechanisms within the model. First, the attention matrix depends on predefined boost invariant pairwise quantities, reducing the network parameters significantly from the original particle transformer models. Second, IAFormer incorporates the sparse attention mechanism by utilizing the "differential attention", so that it can dynamically prioritize relevant particle tokens while reducing computational overhead associated with less informative ones. This approach significantly lowers the model complexity without compromising performance. Despite being computationally efficient by more than an order of magnitude than the Particle Transformer network, IAFormer achieves state-of-the-art performance in classification tasks on the top and quark-gluon datasets. Furthermore, we employ AI interpretability techniques, verifying that the model effectively captures physically meaningful information layer by layer through its sparse attention mechanism, building an efficient network output that is resistant to statistical fluctuations. IAFormer highlights the need for sparse attention in Transformer analysis to reduce the network size while improving its performance.
@@ 1 Introduction
Jets are fundamental objects for LHC collision data, and choosing an appropriate jet tagging algorithm is essential to uncover intriguing physics phenomena. For scenarios involving the production of a high momentum, hadronically decaying heavy particle, such as W and Z bosons or top quarks, a large radius jet is often used to capture entire particles arising from the heavy particle decay.
The recognition of fat jets with the multi-prong structure was first noted in Ref. [19], where jet substructure techniques emerged as a powerful method for suppressing QCD backgrounds, thereby increasing the sensitivity of Beyond Standard Model (BSM) scenarios at the LHC. The pronged structure associated with different heavy resonant particles is dictated by the final state particles of their decays. Over time, jet substructure methodologies have evolved, leading to Deep Learning (DL) based classification techniques for tagging heavy particles [6, 17, 51, 60, 18, 20, 11, 49, 28, 76, 27, 40, 4, 64, 45, 3, 8, 38, 41, 43]. These approaches are now widely applied in LHC analyses conducted by ATLAS and CMS [73, 48, 74, 1, 7, 53]. They are also employed in scenarios with multiple heavy particle final states [9, 42, 32] and in extracting fundamental parameters of the theory [25, 21, 54, 75, 15]. In contrast, processes that generate spatially isolated quarks or gluons are typically reconstructed using one or more small radius jets. This is followed by flavour tagging, a technique that identifies the initiating parton, whether an up, down, strange, charm, bottom quark, or a gluon. Bottom and charm hadrons (primarily B and D hadrons) have relatively long lifetimes, allowing them to travel a short distance within the detector before decaying. Their decays produce displaced tracks originating from a secondary vertex, distinct from the primary collision point. By leveraging machine-learning techniques, this vertex displacement information, combined with subsequent decay properties, significantly improves tagging efficiency [5, 72, 37, 39]. Additionally, studies have explored the potential for strange-quark identification using ML techniques [31, 65, 30]. Since gluons belong to the octet representation of QCD SU(3) symmetry, their jet characteristics differ from those of quarks, which are in the triplet representation. Consequently, ML-based methods have also been developed to distinguish gluon-initiated jets from quark-initiated ones [56, 22, 2, 77, 82, 36].
@@
Recently, Transformer networks have demonstrated significant improvements in jet tagging tasks[69, 79, 43]. Standard Transformers process jet constituents along with their corresponding features, which include kinematic information for each hadron inside the jet. In this setup, the network assigns weights to each particle based on its feature importance in the overall classification decision.
Incorporating features derived from particle-pair interactions further enhances network performance, as these features reflect the radiation patterns inside the jet [28]. The pairwise interactions have been integrated into the attention mechanism either using an interaction matrix as a bias for the computed attention scores [69] or replacing the Query and Key matrices with the interaction matrix itself [79]. In both approaches, the number of attention heads must be equal to or larger than the number of features in the interaction matrix. Additionally, the interaction matrix remains fixed across attention heads and is injected into each attention layer without updates. These rigid structure requirements can disrupt the learning of attention patterns across different layers, potentially limiting the overall effectiveness. IAFormer addresses these challenges by allowing the interaction matrix to be updated across different attention heads and Transformer layers. Furthermore, the interaction matrix is refined through skip connections and propagated to later attention layers.
@@
IAFormer architecture demonstrates superior performance compared to previously introduced Transformer-based models when evaluated on top tagging and quark-gluon tagging tasks with a significantly reduced network size. Moreover, IAFormer is designed as a general framework that can be adapted for various classification tasks. The implementation is publicly available at IAFormer GitHub, and detailed usage instructions are provided in the appendix.
This paper is organized as follows: Section 2 discusses various Transformer network architectures, including the structure of IAFormer. Section 3 presents validation results, comparing IAFormer with existing networks for top tagging and quark-gluon tagging tasks. In Section 4, we employ different interpretability methods to analyze the learned attention patterns in IAFormer during the top tagging task. Section 5 provides a discussion and conclusion.
@@ 2 Transformers for jet tagging
Transformers were originally introduced for natural language processing, particularly for sequence-to-sequence tasks like machine translation [78], which were later modified for collider physics. The core of Transformer is the self-attention mechanism, which dynamically assigns importance to different input components. One of the most significant advantages of self-attention is its ability to capture long-range dependencies, as it considers all elements of the input simultaneously rather than relying on sequential processing. Moreover, modifications to the self-attention formula by incorporating the particle pair interactions enhance the network’s performance.
In jet tagging tasks, the outcome of an event should remain unchanged regardless of the order in which particles are processed. Therefore, the designed DL networks should treat particles as an unordered set, commonly referred to as a "particle cloud". By representing hardons inside the jet in a permutation-invariant manner, particle cloud models overcome the challenges posed by combinatorial ambiguities. The particle-based methods also provide the flexibility to incorporate additional particle attributes, such as particle identification and vertex information.
Several machine learning architectures have been developed for jet tagging under the permutation invariant framework, including Deep Sets [57], Edge Convolution Networks [68], and Transformers [69, 42, 50, 14, 45, 41]. Deep Sets can achieve state-of-the-art performance but require a large latent space, making them computationally demanding. Edge Convolution Neural Networks (EDGCNN) address this challenge by leveraging local neighborhood information within the particle cloud to enhance learning. For the case of particle transformers, attention weights are computed for every particle interaction within the dataset, and outputs are aggregated to keep the permutation invariance, so that the network output remains unaffected by the order of the input. However, self-attention comes with computational challenges, particularly its quadratic scaling with input size. Various improvements, such as sparse attention [23], have been proposed to address this issue.
In the following, we first introduce the Plain Transformer and its modified version, including the interaction matrix, then present the new network, IAFormer. IAFormer significantly reduces the computational burden of attention matrix calculation by incorporating prefixed pairwise information with trainable weights. Moreover, it reduces statistical fluctuations by introducing a dynamic spatial attention mechanism, referred to as differential attention.
@@ 2.1 Plain Transformer
The core component of a transformer model is the attention matrix, a weighted sum of the input elements. The weights are determined based on the relevance of each element to the others. For the particle transformer, the attention weights are computed in a permutation-invariant manner, and the resulting model is able to process sets efficiently.
Mathematically, the attention mechanism operates as follows: The input to the multi-head attention $X$ has a grid-like structure with dimensions $I\times J$, with $I$ indicating the number of particles and $J$ indicating the relevant features for each particle in the dataset.
For given $X_{ij}$, the feature vector $X_{*j}$ is embedded into a higher dimension via an MLP. The embedding to a higher dimension helps the network to capture meaningful complex patterns and nonlinear relationships that may not be apparent in the original feature space; therefore, it enhances the model’s ability to distinguish different classes for deep learning tasks like classification and anomaly detection. High-dimensional embeddings also provide richer representations that facilitate downstream tasks such as contrastive learning and clustering. The size of the embedded features is determined by the number of neurons in the final MLP layer, as
$X^{{\rm embed},(l)}_{i,j}=\sigma\left(X^{(l-1)}_{i,j^{\prime}}W^{(}l)_{j^{\prime},j}+b^{(}l)_{j}\right)\,$, (1)
@@
Once the input features are mapped to a higher dimension, the inputs are passed to three separate linear layers, forming the Query, Key, and Value matrices as
$Q^{I\times J}_{i,j}\equiv X^{\rm embed}_{i,j^{\prime}}\cdot W^{Q}_{j^{\prime},j}\,,\hskip 11.38109ptK^{I\times J}_{i,j}\equiv X^{\rm embed}_{i,j^{\prime}}\cdot W^{K}_{j^{\prime},j}\,,\hskip 11.38109ptV^{I\times J}_{i,j}\equiv X^{\rm embed}_{i,j^{\prime}}\cdot W^{V}_{j^{\prime},j}\,$, (2)
where $Q$, $K$, and $V$ are the query, key, and value matrices, respectively, used to compute the attention for the dataset. The suffix $I\times J$ indicates the dimension of the matrices, where $I$ is the number of particle tokens and $J$ is the number of embedded features. The weight matrices have the dimensions of $I\times I$ to preserve the dimensions of the embedded input dataset. The attention score $\alpha$ is defined as:
@@
where suffix $I\times I$ of $\alpha$ indicates the dimension of the attention weight matrix $\alpha$. In the last line, the sum runs over all components $I\times I$, and $d$ is the length of the particle tokens.
@@ 2.2 Transformer with interaction matrix
The pairwise interaction matrix is the matrix of the prefixed quantities calculated from the pair of particle tokens. For the jet classification, the precomputed variables are mass, relative angles, $k_{T}$, etc, of the pair of jet constituents. The standard Transformer architecture does not inherently support the integration of the pairwise interaction matrix within its attention mechanism. The pairwise interaction matrix has dimensions of (particles×particles×pairwise features)($\text$$\times\text{particles}\times\text${pairwise features}). A straightforward approach to incorporating the pairwise interaction matrix is by embedding it directly into the attention score, equation 3. The feature dimension is adjusted to match the number of attention heads using a linear embedding layer. Consequently, the attention score structure for a single attention head, incorporating the interaction matrix, takes the form that is utilized by ParT[69],
@@
with ${\mathcal{I}}_{i,j}$ is the pairwise interaction matrix for a single head after the embedding. In this structure, the interaction matrix acts as a bias in the attention score.
While this approach effectively integrates the interaction matrix into the attention score, it has certain limitation. The feature dimension of the interaction matrix needs to be reshaped to be compatible with the specified number of attention heads. This occurs because $Q$ and $K$ must accommodate the same number of attention heads in this setup.
@@
While this approach reduces network complexity and improves performance, it has a key limitation. Because the interaction matrix is added independently at each Transformer layer, a mismatch can arise between the feature representations of the pairwise interaction matrix and the input $X_{i,j}$. Since the attention mechanism assigns weights to all particle tokens in the dataset, this misalignment can lead to attention being allocated to irrelevant particles, such as soft hadrons.
@@ 2.3 IAFormer
IAFormer is a transformer-based architecture, but significant modifications have been made compared to the ParT setup, as illustrated in Figure 1. Instead of computing the attention score through the multiplication of the Query and Key matrices, IAFormer applies the softmax function directly to a trainable interaction matrix of the form $W\mathcal{I}$. This interaction matrix is independently optimized for each attention head within the individual attention layers, enabling the network to dynamically learn discriminative patterns among different jets with effectively reduced size.
@@
Additionally, a dynamic sparse attention pattern is included following the idea of “differential attention” [80]. The differential attention utilizes the attention scores as the difference between two separate softmax attention maps. The subtraction cancels noise, promoting the sparse attention patterns dynamically. The model has a trainable parameter $\beta$ to control the level of the subtraction. The parameter is initialized randomly and shared across all attention heads within each attention layer. This mechanism suppresses attention scores for less relevant particle tokens, enhancing the model’s focus on important interactions.
@@ 2.3.1 Dynamic sparse attention via differential attention
Sparse attention is a technique used in Transformer models to reduce the computational complexity of the self-attention mechanism by limiting the number of key-query pairs attending to the classification. Instead of computing attention over all tokens in a sequence, sparse attention selectively attends to a subset of tokens based on predefined patterns, learned structures, or efficient approximations. This reduces the memory and computational cost from quadratic to linear or sub-quadratic complexity, making it possible to process longer sequences efficiently.
Different types of sparse attention have been introduced, such as fixed pattern and learnable sparse attention. Fixed pattern sparse attention is a type of sparse attention mechanism where each token attends only to a predefined subset of tokens based on a fixed rule [33, 66, 10]. On the other hand, Learnable sparse attention dynamically determines which tokens to attend to others based on learned patterns. Learnable sparse attention allows the model to adaptively focus on the most relevant tokens in different contexts, leading to improved efficiency. In this paper, we utilize a dynamic sparse attention mechanism incorporated in the transformer called “differential attention” which was first introduced in [80],
@@
with $\beta$ is a learnable vector and $W_{1}$$W_{2}$ are two learnable matrices. During the training process, the value of $\beta$ vector is optimized to demote the attention score assigned to less relevant tokens, resulting in the attention score matrix to be sparse. This results in an implicit sparsity, where tokens gradually focus only on meaningful interactions while ignoring less relevant ones. The self-attention mechanism is guided by this dynamic sparsity, reducing computational overhead while preserving essential long-range dependencies. The IAFormer shows better performance compared with ParT using a notably smaller number of parameters.